Question 9 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price in both countries. The table below shows the combinations of the two goods that each country can produce in one year using the same amounts of capital and labor. What is the opportunity cost of producing one ton of grapes (in terms of ships) for France? Provide your answer as a number rounded to two decimal places. Do not include any symbols, such as S -%, or , in your answer. Spain France Ships Grapes (in tons) 75 Grapes (in tons) 50 40 30 20 10 Ships 10 20 30 40 50 45 30 15 30 45 60 75 Your Answer: Answer Question 10 (1 point) ish are sold for the same price in
Question 10 (1 point) Suppose that Spain and France both produce ships and grapes, which are sold for the same price in both countries. The table below shows the combinations of the two goods that each country can produce in one year using the same amounts of capital and labor. What is the opportunity cost of producing one ship (in terms of tons of grapes) for Spain? Provide your answer as a number rounded to two decimal places. Do not include any symbols, such as S, , %, or , in your answer. France Ships Grapes (in tons) 75 60 45 30 15 Ships 10 20 30 40 50 15 30 45 60 75 Grapes (in tons) 50 40 30 20 10 Your Answer: Answer
Answer #1

9) opportunity cost of grapes for France= 75/50= 1.5 ships

Answer=1.5

10) opportunity cost of one ship in Spain= 75/50= 1.5 tons of grapes

Answers- 1.5